A gymnasium is 16 m high. By what percent is the air pressure at the floor greater than the air pressure at the ceiling? Assume that pressure at the bottom is 1 atm and the density of air stays constant. The density of air is 1.29 kg/m3

1
Expert's answer
2021-10-02T14:55:33-0400

"\Delta p=\rho g(h_2-h_1)=1.29\cdot9.81\cdot16\approx202\ (Pa)"


"h_1=101\ 325\ (Pa)"


"h_2=101\ 123\ (Pa)"


"e=100-101123\cdot100/101325=0.2\ (\%)" . Answer